What is the answer for the question 2(-1-2n)=22
Answer by ReadingBoosters(3246)

Starting with,
2(-1-2n) = 22
Distribute the 2 on the left
-2 - 4n = 22
-2 + 2 - 4n = 22 + 2
-4n = 24
-4n/-4 = 24 / -4
n = -6